The length of a rectangle is 6 feet less than twice the width. The perimeter is 66 feet. Find the length and width.
Answer by checkley77(12569)

L=2W-6
2L+2W=66
2(2W-6)+2W=66
4W-12+2W=66
6W=66+12
6W=78
W=78/6
W=13 ANS. FOR THE WIDTH.
L=2*13-6
L=26-6
L=20 ANS. FOR THE LENGTH.
PROOF:
2*20+2*13=66
40+26=66
66=66